Solve graphically y = 2x - 6 and y = -2x + 10. Write down the co-ordinates of the point of intersection of the two lines
Answer by mananth(16949)

-2 x + 1 y = -6 .............1
2 x + 1 y = 10 .............2

equation1
when x= 0, y= -6, ( 0 , -6 )
when y = 0, x= 3 ( 3 , 0)
Equation 2
when x= 0, y= 10, ( 0 , 10 )
when y = 0, x= 5, ( 5 , 0)

(4,2) point of intersection
m.ananth@hotmail.ca